Exercise 3-10 Computing and interpreting profit margin LO A1 r answers to 1 decímal
Answer #1

Solution

Profit margin = net income/net sales

Company a: 6298/51200=12.3%

Company b: 101617/455680=22.3%

Company c: 105938/291840=36.3%

Company d: 79872/1664000=4.8%

Company e: 86415/496640=17.4%

According to profit margin ratio company c is the most profitable.